OpenStudy (anonymous):

integral 2x / (x+3) (3x+1) dx

OpenStudy (anonymous):

I know the following steps: 2x / (x+3) (3x+1) dx = A/(x+3) + B/(3x+1) 2x = A(3x+1) + B(x+3) But i don't know how to calculate A and B

OpenStudy (helder_edwin):

first put x=-3 and then x=-1/3